OpenStudy (anonymous):

If the measures of the angles of a triangle are in the ratio 1:3:5 , the number of degrees in the measure of the smallest angle is a. 10 b. 20 c. 60 d. 180

OpenStudy (anonymous):

9a = 180 Divide both sides by 9 a = 20 . Substituting back into the formulas we defined. a = 20 b = 3a = 3(20) = 60 c = 5a = 5(20) = 100 . Check the work. Does 20 + 60 + 100 = 180? Yes.

OpenStudy (anonymous):

where did the 9 come from @kumar2006